OpenStudy (anonymous):

medal will be rewarded algebra 1!

jimthompson5910 (jim_thompson5910):

you have 2 points (0, -4) and (1, -3) use the slope formula m = (y2 - y1)/(x2 - x1) m = (-3 - (-4))/(1-0) m = (-3 + 4)/(1-0) m = 1/1 m = 1

OpenStudy (anonymous):

Ok thanks! It is still a little confusing though....

OpenStudy (campbell_st):

plot the points  the slope is rise over run